OpenStudy (anonymous):

Solve via a method of factoring. Hint: write it in ax^2+bx+c=0 form first. a^2 = 4(3a-9)

OpenStudy (zzr0ck3r):

a^2 = 4(3a-9) a^2 = 12a -36

OpenStudy (zzr0ck3r):

now we get "everything" on one side. a^2 -12a + 36 = 0

OpenStudy (zzr0ck3r):

do you know what to do from here?